OpenStudy (anonymous):

solve: tan x = sqrt(3)

OpenStudy (maheshmeghwal9):

\[\tan x =\sqrt{3} \implies x = \tan^{-1}\sqrt{3}.\]\[\text{Thus ;} \space \space x = 60^o = \frac{\pi}{3}.\]

OpenStudy (anonymous):

thank you !!

OpenStudy (maheshmeghwal9):

:)